OpenStudy (anonymous):

6⋅x2 = 5⋅x + 1 solve using quadratic equation

OpenStudy (anonymous):

\[(-b \pm \sqrt{b^2-4ac} ) / 2a\]\[6x^2 = 5x + 1\] First, subtract (5x+1) from both sides so that the right side is equal to zero. You could subtract 6x^2 from both sides but I prefer the leading term to be positive. \[6x^2 - 5x - 1 = 0\] This is a form of the equation ax+bx+c = 0, so a = 6 b = -5 c = -1 Plug this into the equation above: \[(-b \pm \sqrt{b^2-4ac} ) / 2a\] And you get \([-(-5) \pm \sqrt{(-5)^2 - 4(6)(1)})/2(6)\] Which simplifies to\[(5 \pm \sqrt{25 - 24}) / 12\] or\[(5 \pm \sqrt{1}) / 12\] Which is \[(5 \pm 1)/12\] Because the square root of 1 is 1 (1*1=1) Now, you have to take both the + and - possibilities of the equation:  So, in short, x = 1/2 and x = 1/3